Peter is the owner of a hotel. He has found that when he charges the nightly cost of $124, the average occupancy is 412. In addition, Peter has found that with every $10 increase in the average nightly cost, the hotel occupancy decreases by an average of 15.
Which of the following functions describes the hotel's profit in terms of x, the number of $10 increases over $124?
A. P(x) = $51,088 - $1,448x - $150x^2
B. P(x) = $51,088 + $5,980x + $150x^2
C. P(x) = $51,088 + $2,260x - $150x^2
D. P(x) = $51,088 - $1,960x - $150x^2

Answer by stanbon(75887)

Peter is the owner of a hotel. He has found that when he charges the nightly cost of $124, the average occupancy is 412. In addition, Peter has found that with every $10 increase in the average nightly cost, the hotel occupancy decreases by an average of 15.
Which of the following functions describes the hotel's profit in terms of x, the number of $10 increases over $124?
Revenue = (unit price)*(units rented)
P(x) = (124+10x)(412-15x) = 51,088 + 2260 - 150x^2
-------------------------------------------------------
Cheers,
Stan H.
-----------------
A. P(x) = $51,088 - $1,448x - $150x^2
B. P(x) = $51,088 + $5,980x + $150x^2
C. P(x) = $51,088 + $2,260x - $150x^2
D. P(x) = $51,088 - $1,960x - $150x^2